OpenStudy (anonymous):

Need Help Please If sin y° = s divided by 8. and tan y° = s divided by t., what is the value of cos y°? cos y° = 8s cos y° = 8t cos y°=t divided by 8. cos y°=8 divided by t.

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@dmezzullo

OpenStudy (anonymous):

its t divided by 8

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

jw

OpenStudy (anonymous):

im not very good at trig functions

OpenStudy (anonymous):

cos= sin/tan (s/8)/(s/t) (s/8)*(t/s) st/8s cos y = t/8

OpenStudy (anonymous):

that help?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

medal me?

OpenStudy (anonymous):

i gave you one

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

could you help me with another?

OpenStudy (anonymous):

I can try but i'm not to good at them either. but i'll give it my best shot and see what I can do.